EDITORIAL

Terry Goodman

As another school year comes to a close there are several things that we would like to bring to the attention of our readers and those who submit manuscripts for publication.

Manuscripts submitted for publication should be no more than ten typed pages in length. While some exceptions to this guideline may, at times, be necessary, we will, in general, adhere to it in order to keep each issue of MJMS within the 50-60 page range.

Overall, the number of manuscripts we are receiving is adequate, there are some areas in which we would like to have more articles. We encourage the submission of manuscripts dealing with topics in number theory, combinatorics, statistics, computer science, and mathematics education.

We continue to encourage the submission of problems for our Problems section. Many of our readers indicate that this section of the journal is particularly interesting and useful. Problems for student solutions only are most welcome also.

Finally, we are looking for individuals who could serve as referees for manuscripts dealing with applied mathematics topics. If you are interested, please complete the Referee Application form on page 113.